OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 2x + 3y = 4 x - 2y = -5 A.) 2 B.) -2 C.) 1 D.) -1

OpenStudy (stamp):

2x + 3y = 4 x - 2y = 5 2x + 3y = 4 2x - 4y = 10 7y = -6 y = -6/7 Use this y value to solve for your x, you can substitute the y value back in to either equation to solve for x.

OpenStudy (anonymous):

I really don get how u did that

OpenStudy (stamp):

@dusty Go find an example in your textbook, read it, then read my work.

OpenStudy (anonymous):

i dont have a txt book lol i do my stuff online

OpenStudy (anonymous):

i do my stuff online too but ill give you an explanation

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

@summer1980 how did you get that answer????

OpenStudy (anonymous):

multiply the first equation by 2 and the second one by 3 then add them you'll get 7x=7